OpenStudy (amonoconnor):

Would someone be willing to check my work for this one? "Find [d^2f/dx^2](x) if f(x) = log(base2)x." Any and all help is greatly appreciated!

OpenStudy (irishboy123):

i'd switch it out of base 2 before even starting...

OpenStudy (amonoconnor):

Here's what I did: \[f(x) = \log_{2}x \] \[f'(x) = \frac{1}{x* \ln 2}\] \[f''(x) = \frac{(0)*(xln2) - 1*[(1)*\ln2 + x(1/2)]}{(xln2)^2}\]

OpenStudy (amonoconnor):

I used the formula for the derivative of a log with an arbitrary base for f'(x), and Quotient Rule for f''(x).

OpenStudy (amonoconnor):

Is this the correct final answer? \[f''(x) = \frac{-\frac{1}{2}x - \ln2}{(xln2)^2}\]

OpenStudy (irishboy123):

no, it's much more straightforward than that

OpenStudy (amonoconnor):

:/ Rgg... Would you walk me through it?

OpenStudy (irishboy123):

\(f'(x) = \dfrac{1}{x* \ln 2} = \dfrac{1}{\ln 2} . \dfrac{1}{x}\) what you want for the second deriv is \((\dfrac{1}{x})'\)

OpenStudy (irishboy123):

\[\dfrac{1}{x} = x^{-1}\] you can use that rule \((x^n)' = n x^{n-1}\)

OpenStudy (amonoconnor):

Then what happens to the 1/ln2?

OpenStudy (irishboy123):

treat \(\dfrac{1}{\ln 2} \) as a constant because it is a constant...

OpenStudy (amonoconnor):

So it's like using the Product Rule, and 1/ln2 just stays there?

OpenStudy (irishboy123):

if you want to see it that way but i think that is still overcomplicating it... think: \( (Ax^{n})' = n A x^{n-1} \)

OpenStudy (irishboy123):

just as \((A \sin x)' = A \cos x\) \( (A e^{x})' = A e^x \) \((A \ln x)' = \frac{A}{x}\) constant just hangs there....but yes you could use a product rule if you wanted too but it would be unnecessary....

OpenStudy (amonoconnor):

Is this the final answer then? \[f''(x) = \frac{1}{\ln 2}*(-1x^{-2})\] \[= \frac{1}{\ln2}*\frac{-1}{x^{-2}}\] \[= \frac{-1}{(x^{-2})\ln2}\] ?

OpenStudy (amonoconnor):

pellet, I see a mistake. Take the negative from the x^-2 after I put it in the denominator!! My bad!

OpenStudy (amonoconnor):

Thank you:)

OpenStudy (irishboy123):

mp constants flow through like this... \(\large f'(a)=\lim_{h\to 0}\frac{f(a+h)-f(a)}{h}\) \(\large A f'(a)=\lim_{h\to 0}\frac{Af(a+h)-Af(a)}{h}\)

OpenStudy (amonoconnor):

I know how constants work, I just didn't know that ln(a) would be a constant, and not have to have its derivative taken again. :)

OpenStudy (irishboy123):

great!